X-y=-3
4x+y=-2
How do I solve this and graph it?

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

x-y=3
y=x-3
-------------
4x+y=-2
y=-4x-2
-------------
4x+x-3=-2
5x=1
x=1/5
or
x=0.2
-----------
4*0.2+y=-2
0.8+y=-2
y=-2-0.8
y=-2.8
-----------
(0.2,2.8)intersection

Answer by stanbon(75887)

X-y=-3
4x+y=-2
How do I solve this and graph it?
==============
Solve each for "y"::
y = x+3
y = -4x-2
----------
Solve for "x"::
x+3 = -4x-2
5x = -5
x = -1
--
Then y = x+3 = 2
Solution:: (-1,2)
-------------------------
-----------
Cheers,
Stan H.
------------